OpenStudy (trojanpoem):

Find the limit :

OpenStudy (trojanpoem):

\[\lim_{x \rightarrow \infty} x[ \sin( \frac{ a }{ x }) - \tan(\frac{ a }{ x })]\]

OpenStudy (turingtest):

\[\lim_{x \rightarrow \infty} x[ \sin( \frac{ a }{ x }) - \tan(\frac{ a }{ x })]\\\lim_{t \rightarrow 0} \frac1t[ \sin( at) - \tan(at)]\]distribute and you should find some familiar limits

OpenStudy (trojanpoem):

You assumed x = 1/t sin(at)/t - tan(at)/t = a -a = 0 ?

OpenStudy (turingtest):

seems that way

OpenStudy (trojanpoem):

Thanks so much, May you help me with another one ?

OpenStudy (turingtest):

sure :D

OpenStudy (turingtest):

but you should post separately

OpenStudy (trojanpoem):

I will and tag your name.